OpenStudy (anonymous):

what is the standard form of the slope -1/3?

OpenStudy (anonymous):

would it be x+3y=3?...

OpenStudy (anonymous):

no

OpenStudy (anonymous):

we know that the standard form is y=mx+b

OpenStudy (anonymous):

x+3y=8 would be your correct answer